I need to convert a paint that I bought that has 7% of the dye color I want but the shade I like best calls for only 3% dye color. How much pure white needs to be mixed with the 7% to get .7 gallons of the 3%dye paint.
Answer by mananth(16946)

percent ----------------quantity
White paint 0----------------x gallons
Paint with dye 7----------------0.7-x gallons
Total 3.00% ----------------0.7 gallons

0x+7(0.7-x)=3*0.7
4.9-7x=2.1
-7x=2.1-4.9
-7x=-2.8
/-7
x=0.4 gallons white paint
0.3 gallons 7.00%